# PDMLP: Patch-based Decomposed MLP for Long-Term Time Series Forecasting

### v1

@@ 1 Introduction
Figure 1: The self-attention scores of a 2-layer Transformer with different Patch size trained on ETTh1. We follow the setup of PathcTST Nie et al. (2022), retaining only the Encoder while replacing the Decoder with a simple MLP, and using a channel independent approach. A patch size of 1 is equivalent to the original Transformer, indicating that time series data often exhibits a trend of being segmented into patches Zhang & Yan (2022); Tang & Zhang (2023), and an increase in patch size can mitigate this to some extent.
@@
  • Does the impressive performance of the Patch-based Transformer derive from the inherent strength of the Transformer architecture, or is it merely due to the use of patches as the input representation?
@@
In this paper, we specifically discuss the three issues mentioned above and introduce the PDMLP, a novel and concise Patch-based decomposed MLP model tailored for LTSF task. PDMLP is entirely based on fully connected networks and incorporates the concept of patche. Specifically, we employ a patch methodology to embed the original sequences into representational spaces, then extract the smooth components and noisy residuals of the series using a simple moving average technique for separate processing. We independently process stochastic noise across channels and facilitate semantic information interchange between variables through channel mixing. Upon evaluation on numerous real-world datasets, the PDMLP model demonstrated state-of-the-art (SOTA) performance. To be precise, our contributions can be summarized in three aspects:
  • We analyze the effectiveness of Patches in time series forecasting and propose a Multi-Scale Patch Embedding (MPE) approach. Unlike previous embedding methods that used a single linear layer, MPE is capable of capturing multiscale relationships between input sequences.
@@ 2 Patch for Long-Term Time Series Forecasting
So, is a larger patch always better? In Figure  2(b), a larger patch does not seem to achieve better results, which may be partly influenced by the size of the hidden layer ($d_{model}$). With a larger patch size, more time points are projected into a fixed-size $d_{model}$, which may lead to excessive compression. As shown in Figure  2(c), as $d_{model}$ increases, the performance of the model gradually improves, indicating that larger patches perform better. In the extreme case, a patch that treats the entire input sequence as one patch and projects it into a vector is similar to iTransformer Liu et al. (2023). However, a larger $d_{model}$ indicates that there are more parameters for the model to learn, which can easily lead to underfitting and result in decreased model performance.
@@ 3.1 Multi-Scale Patch Embedding
Time series analysis relies on the accurate identification of local information within the sequence and optimization of the model. The use of patches can provide the model with a local view of the time series in the short term, thereby enhancing the representation of local information within the sequence and enabling the model to learn and capture these local features more precisely. However, traditional methods typically employ patches of a single-scale to embed the raw time series, which makes the model inclined to learn single-scale temporal relationships while neglecting the multi-scale nature and complexity of time series data.
@@
To decompose the multivariate time series $\mathcal{X}$ into a univariate series $x$, we have developed a suite of patches $\mathcal{P}$ across various scales to process $x$. For a particular scale $p\in\mathcal{P}$, $x$ is first divided into non-overlapping patches where the length of patch corresponds to $p$, thus the patching process yields the sequence of patches $x_{p}\in\mathbb{R}^{N\times p}$, where $N$ is the number of patches. Subsequently, we employ a single-layer linear layer to embed the patches $x_{p}$, resulting in latent vectors $x_{e}\in\mathbb{R}^{N\times d}$, $d$ is the embedding dimension, the $d$ corresponding to different scales patches can be different. These latent vectors are then unfolded and concatenated with latent vectors obtained from patches at alternative scales, thus assembling the ultimate embedding vector $X\in\mathbb{R}^{1\times d_{model}}$, $d_{model}$ is the final embedding dimension input into the model. This multi-scale patch strategy permits the model to capture and learn the intrinsic dynamics of time series at different levels, thereby providing more precise insights when forecasting future trends and patterns.
@@ A.1 Patch-Based Model
Patch-based methods have already demonstrated their immense potential across various data modalities. In the field of Computer Vision (CV), traditional Convolutional Neural Networks (CNNs) naturally segment images into patches and process each patch to learn local features Krizhevsky et al. [2012]. Vision Transformer (ViT) Dosovitskiy et al. [2020] was the first to integrate the concept of patche into the Transformer Vaswani et al. [2017] model, outperforming CNN-based approaches on large datasets and breaking the inductive bias associated with CNNs. Subsequent models such as BEiT Huang et al. [2024] and MAE He et al. [2022] also segment images into patches as input for the Transformer.
In the realm of time series analysis, PatchTST Nie et al. [2022] innovatively segments time series into multiple patches to extract localized semantic information for subsequent Transformer processing. MTSMAE Tang & Zhang [2022] introduced the masked autoencoder method to self-supervised pre-training for time series, learning deep semantic associations through masked patches. iTransformer Liu et al. [2023] further takes a reverse approach by treating an entire time series as a single patch for processing. Additionally, patch methods have been employed within CNN models, such as PatchMixer Gong et al. [2023]. To address the potential loss of temporal information by permutation-invariant self-attention mechanisms, it incorporates a novel permutational variant convolutional structure to preserve temporal information. Relying entirely on depthwise separable convolutions, PatchMixer extracts both local features and global correlations within a single-scale architecture.
Furthermore, TSMixer Ekambaram et al. [2023] draws inspiration from the success of the MLP-Mixer Tolstikhin et al. [2021] model in CV, combining patches with MLP to propose a lightweight, all-MLP neural architecture specifically designed for MTS and representational learning of patch time series. Concurrently, HDMixer aims to capture semantically enriched patches through effective modeling of hierarchical interactions, designing a Length-Extendable Patch adaptable for MTS, enriching the boundary information of patches and alleviating semantic incoherence within sequences.
@@ Appendix D Attention score map of Transformer
The traditional Transformer does not adopt the idea of patche but instead directly embeds the time series at the point-level. Much research indicates that this approach tends to lead to the Transformer processing a large amount of redundant features Tang & Zhang [2022], resulting in lower attention values for certain time points, exhibiting a long-tail distribution with attention scores clearly showing a form of blockiness Zhou et al. [2021]. Figure  12 shows a typical attention score map of the original Transformer for time series forecasting. We selected three datasets of different sizes based on different sampling scales and trained a two-layer Transformer with eight heads, ultimately selecting the attention scores from the second layer. We can see that attention values tend to patch, i.e., close data points have similar attention weights. This suggests that directly computing attention on time points is not advisable for Transformers. However, as shown in Figure  1, creating patches from the input time series can only mitigate this blockiness and does not completely resolve it.

### v2

@@ 1 Introduction
Figure 1: The self-attention scores of a 2-layer Transformer with different Patch size trained on ETTh1. We follow the setup of PathcTST Nie et al. (2022), retaining only the Encoder while replacing the Decoder with a simple MLP, and using a channel independent approach. A patch size of 1 is equivalent to the original Transformer, indicating that time series data often exhibits a trend of being segmented into patch Zhang & Yan (2022); Tang & Zhang (2023), and an increase in patch size can mitigate this to some extent.
@@
  • Does the impressive performance of the Patch-based Transformer derive from the inherent strength of the Transformer architecture, or is it merely due to the use of patch as the input representation?
@@
In this paper, we specifically discuss the three issues mentioned above and introduce the PDMLP, a novel and concise Patch-based decomposed MLP model tailored for LTSF task. PDMLP is entirely based on fully connected networks and incorporates the concept of patch. Specifically, we employ a patch methodology to embed the original sequences into representational spaces, then extract the smooth components and noisy residuals of the series using a simple moving average technique for separate processing. We independently process stochastic noise across channels and facilitate semantic information interchange between variables through channel mixing. Upon evaluation on numerous real-world datasets, the PDMLP model demonstrated state-of-the-art (SOTA) performance. To be precise, our contributions can be summarized in three aspects:
  • We analyze the effectiveness of patch in time series forecasting and propose a Multi-Scale Patch Embedding (MPE) approach. Unlike previous embedding methods that used a single linear layer, MPE is capable of capturing multiscale relationships between input sequences.
@@ 2 Patch for Long-Term Time Series Forecasting
So, is a larger patch always better? In Figure  2(b), a larger patch does not seem to achieve better results, which may be partly influenced by the size of the hidden layer ($d_{model}$). With a larger patch size, more time points are projected into a fixed-size $d_{model}$, which may lead to excessive compression. As shown in Figure  2(c), as $d_{model}$ increases, the performance of the model gradually improves, indicating that larger patch perform better. In the extreme case, a patch that treats the entire input sequence as one patch and projects it into a vector is similar to iTransformer Liu et al. (2023). However, a larger $d_{model}$ indicates that there are more parameters for the model to learn, which can easily lead to underfitting and result in decreased model performance.
@@ 3.1 Multi-Scale Patch Embedding
Time series analysis relies on the accurate identification of local information within the sequence and optimization of the model. The use of patch can provide the model with a local view of the time series in the short term, thereby enhancing the representation of local information within the sequence and enabling the model to learn and capture these local features more precisely. However, traditional methods typically employ patch of a single-scale to embed the raw time series, which makes the model inclined to learn single-scale temporal relationships while neglecting the multi-scale nature and complexity of time series data.
@@
To decompose the multivariate time series $\mathcal{X}$ into a univariate series $x$, we have developed a suite of patch $\mathcal{P}$ across various scales to process $x$. For a particular scale $p\in\mathcal{P}$, $x$ is first divided into non-overlapping patch where the length of patch corresponds to $p$, thus the patching process yields the sequence of patch $x_{p}\in\mathbb{R}^{N\times p}$, where $N$ is the number of patches. Subsequently, we employ a single-layer linear layer to embed the patch $x_{p}$, resulting in latent vectors $x_{e}\in\mathbb{R}^{N\times d}$, $d$ is the embedding dimension, the $d$ corresponding to different scales patch can be different. These latent vectors are then unfolded and concatenated with latent vectors obtained from patch at alternative scales, thus assembling the ultimate embedding vector $X\in\mathbb{R}^{1\times d_{model}}$, $d_{model}$ is the final embedding dimension input into the model. This multi-scale patch strategy permits the model to capture and learn the intrinsic dynamics of time series at different levels, thereby providing more precise insights when forecasting future trends and patterns.
@@ A.1 Patch-Based Model
Patch-based methods have already demonstrated their immense potential across various data modalities. In the field of Computer Vision (CV), traditional Convolutional Neural Networks (CNNs) naturally segment images into patch and process each patch to learn local features Krizhevsky et al. [2012]. Vision Transformer (ViT) Dosovitskiy et al. [2020] was the first to integrate the concept of patch into the Transformer Vaswani et al. [2017] model, outperforming CNN-based approaches on large datasets and breaking the inductive bias associated with CNNs. Subsequent models such as BEiT Huang et al. [2024] and MAE He et al. [2022] also segment images into patch as input for the Transformer.
In the realm of time series analysis, PatchTST Nie et al. [2022] innovatively segments time series into multiple patch to extract localized semantic information for subsequent Transformer processing. MTSMAE Tang & Zhang [2022] introduced the masked autoencoder method to self-supervised pre-training for time series, learning deep semantic associations through masked patches. iTransformer Liu et al. [2023] further takes a reverse approach by treating an entire time series as a single patch for processing. Additionally, patch methods have been employed within CNN models, such as PatchMixer Gong et al. [2023]. To address the potential loss of temporal information by permutation-invariant self-attention mechanisms, it incorporates a novel permutational variant convolutional structure to preserve temporal information. Relying entirely on depthwise separable convolutions, PatchMixer extracts both local features and global correlations within a single-scale architecture.
Furthermore, TSMixer Ekambaram et al. [2023] draws inspiration from the success of the MLP-Mixer Tolstikhin et al. [2021] model in CV, combining patch with MLP to propose a lightweight, all-MLP neural architecture specifically designed for MTS and representational learning of patch time series. Concurrently, HDMixer aims to capture semantically enriched patch through effective modeling of hierarchical interactions, designing a Length-Extendable Patch adaptable for MTS, enriching the boundary information of patch and alleviating semantic incoherence within sequences.
@@ Appendix D Attention score map of Transformer
The traditional Transformer does not adopt the idea of patch but instead directly embeds the time series at the point-level. Much research indicates that this approach tends to lead to the Transformer processing a large amount of redundant features Tang & Zhang [2022], resulting in lower attention values for certain time points, exhibiting a long-tail distribution with attention scores clearly showing a form of blockiness Zhou et al. [2021]. Figure  12 shows a typical attention score map of the original Transformer for time series forecasting. We selected three datasets of different sizes based on different sampling scales and trained a two-layer Transformer with eight heads, ultimately selecting the attention scores from the second layer. We can see that attention values tend to patch, i.e., close data points have similar attention weights. This suggests that directly computing attention on time points is not advisable for Transformers. However, as shown in Figure  1, creating patch from the input time series can only mitigate this blockiness and does not completely resolve it.